OpenStudy (anonymous):

Solve the equation by completing the square x^2-3x-3=0

OpenStudy (anonymous):

x2 + -3x + -3 = 0 Reorder the terms: -3 + -3x + x2 = 0 Move all terms containing X to the left, all other terms to the right. Add '3' to each side of the equation. -3 + -3x + 3 + x2 = 0 + 3 Reorder the terms: -3 + 3 + -3x + x2 = 0 + 3 Combine like terms: try it from here :)